# Quantum Natural Gradient with Geodesic Corrections for Small Shallow Quantum Circuits

### v2

@@ 1 Introduction
The Quantum Natural Gradient (QNG) algorithm [6], a generalization of the natural gradient [7], is an advanced optimization technique that enhances the optimization process by incorporating the geometry of the parameter space into the update rules. Unlike standard gradient descent, which assumes a flat parameter space, QNG leverages a Riemannian metric defined by the Fubini-Study metric or more generally the quantum Fisher information [8], capturing the infinitesimal changes in distances between quantum states and aligning parameter updates with the natural curvature of the quantum state manifold. A key factor in QNG’s superior performance, as shown in [9], is its ability to identify regions of high negative curvature early in the optimization process, significantly accelerating convergence. These regions play a crucial role in guiding the optimization along paths that lead to faster descent.
@@
This manuscript presents a comprehensive approach to incorporating geodesic corrections into the QNG for applications in variational quantum algorithms. In Section 2, we provide an overview of differential geometry and geodesic equations, establishing the foundational mathematical context. Section 3 delves into optimizing idealized variational quantum circuits using Quantum Natural Gradient Descent (QNG), emphasizing its effectiveness in quantum optimization. Section 4 introduces higher-order integrators and derives the update rule for the Quantum Natural Gradient with Geodesic Correction (QNGGC). As the update rule relies on the Christoffel symbols of the second kind, Section 5 is dedicated to efficiently computing these symbols using the parameter-shift rule, enabling direct measurements from quantum circuits. In Section 6, we apply the derived update rule to various examples: Examples 1 and 2 involve analytical calculations for numerical simulations, while Example 3 utilizes quantum software, specifically Qiskit [30], for practical simulations. Finally, in Section 7, we provide an outlook on potential extensions and future research directions, highlighting the integration of geodesic corrections into quantum optimization frameworks.
2 Differential Geometry and Geodesic Equations
@@
A manifold is a topological space that locally resembles Euclidean space and supports a consistent coordinate system. More formally, an $n$-dimensional manifold $\mathcal{M}$ is a set equipped with a collection of coordinate charts $\{(U_{i},\varphi_{i})\}$, where each $U_{i}\subset\mathcal{M}$ is an open subset, and $\varphi_{i}:U_{i}\rightarrow\mathbb{R}^{n}$ is a homeomorphism, meaning that $U_{i}$ is locally similar to $\mathbb{R}^{n}$. For $\mathcal{M}$ to be a differentiable manifold, the transition maps between overlapping charts, $\varphi_{j}\circ\varphi_{i}^{-1}:\varphi_{i}(U_{i}\cap U_{j})\rightarrow\varphi_{j}(U_{i}\cap U_{j})$, must be smooth (infinitely differentiable). This smooth structure allows us to perform calculus on the manifold.
@@
where $g_{ij}=g_{ji}$. The metric allows for measuring lengths and angles on the manifold. The length of a smooth curve $\gamma:[a,b]\rightarrow\mathcal{M}$ is given by:
@@
The Christoffel symbols define the covariant derivative, which maps tensor fields to other tensor fields, adapting to the curvature of the manifold.
The exponential map is a tool in Riemannian geometry that relates the tangent space at a point to the manifold. Given a tangent vector $v\in T_{p}\mathcal{M}$, the exponential map $\exp_{p}$ maps $v$ to a point on the manifold reached by traveling along the geodesic starting at $p$ with initial velocity $v$. Formally,
@@
where $\gamma(\tau)$ is the geodesic satisfying the initial conditions $\gamma(0)=p$ and $\dot{\gamma}(0)=\left.\frac{d\gamma(\tau)}{d\tau}\right|_{\tau=0}=v$. By rescaling the parameter $v$ by a factor of $h$, the following relation holds:
$\text{Exp}_{p}(hv)=\gamma(h).$ (6)
To approximate $\text{Exp}_{p}(hv)$ locally, we expand $\gamma(h)$ for small $h$ using a Taylor series:
$\text{Exp}_{p}(hv)\approx\gamma(0)+h\dot{\gamma}(0)+\frac{h^{2}}{2}\ddot{\gamma}(0)+\mathcal{O}(h^{3}).$ (7)
This expansion provides insight into the behavior of the geodesic near the starting point $p$. The first-order term, $\gamma(0)+h\dot{\gamma}(0)$, corresponds to the initial position and velocity, providing a linear approximation of the geodesic. The second-order term, $\frac{h^{2}}{2}\ddot{\gamma}(0)$, accounts for curvature effects, refining the approximation by incorporating the geodesic’s acceleration. Since $\ddot{\gamma}(0)$ satisfies the geodesic equation, it inherently reflects the manifold’s curvature, capturing both the initial direction and the curvature of the geodesic path.
@@
3 Optimizing Idealized Variational Quantum Circuits with Quantum Natural Gradient
@@
where $V_{j}$ are fixed unitary operators, and $U_{j}(\theta_{j})$ are parameterized gates. Each $U_{j}(\theta_{j})$ is defined as:
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta\,\partial_{j}\mathcal{L}(\bm{\theta}_{t}),$ (11)
where $\partial_{j}:=\frac{\partial}{\partial\theta_{j}}$. Equation (11) can be interpreted as an approximation of the solution to an ordinary differential equation (ODE) using the Euler method:
$\dot{\bm{\theta}}$ $=-\lambda\,\partial_{j}\mathcal{L}(\bm{\theta}),$ (12)
where $\eta=h\lambda$ is the learning rate, with $\lambda$ being a time scale constant that affects the speed but not the trajectory of the system, and $h$ is the step size.
However, this ODE is not invariant under reparameterizations of the parameters $\bm{\theta}$. For instance, if we rescale the parameters $\bm{\theta}\rightarrow 2\bm{\theta}$, the gradient $\partial_{j}\mathcal{L}$ would scale as $\frac{1}{2}\partial_{j}\mathcal{L}$, leading to inconsistencies in the optimization process.
The core of this issue lies in the differential geometric nature of the gradient. The parameter update $\dot{\bm{\theta}}$ transforms as a vector in the tangent space $T_{\bm{\theta}}\mathcal{M}$, while the gradient $\partial_{j}\mathcal{L}$ is a covector (or 1-form) in the cotangent space $T^{*}_{\bm{\theta}}\mathcal{M}$. Since the ODE in Eq. (11) attempts to relate objects in different spaces with distinct transformation rules, it is not an invariant relation.
QNG alleviates this issue by approximately solving an invariant ODE. The key idea is to raise the index of the gradient using a metric tensor $g_{ij}$. By raising the index of $\partial_{j}\mathcal{L}$ on the right-hand side of the gradient descent ODE, the new ODE becomes:
$\dot{\bm{\theta}}$ $=-\lambda g^{ij}\partial_{j}\mathcal{L}(\bm{\theta}),$ (13)
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta g^{ij}\partial_{j}\mathcal{L}(\bm{\theta}_{t}),$ (14)
remains consistent across different parameter spaces. The metric tensor $g_{ij}$ is recognized as the Fubini-Study metric, derived from the real part of the Quantum Geometric Tensor:
@@
4 Higher-order Integrators and Geodesic Correction
@@
$\dot{\gamma}^{i}_{t}(h)$ $=\bm{\theta}_{t+1}=\text{Exp}_{\bm{\theta}_{t}}\left(-h\lambda\,g^{ij}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})\right),$ (16)
@@
$\gamma^{i}_{t}(0)$ $=\bm{\theta}_{t},$ (17)
$\dot{\gamma}^{i}_{t}(0)$ $=-\lambda g^{ij}\partial_{j}\mathcal{L}(\bm{\theta}_{t}).$ (18)
Here, the exponential map Exp translates the current parameter $\bm{\theta}_{t}$ along the geodesic defined by the natural gradient. This approach is effective because it preserves invariance properties under reparameterization, owing to the characteristics of the exponential map. However, directly computing the exponential map is challenging, as it requires solving the geodesic equation exactly. To approximate this computation, we explore using the first and second-order derivatives as detailed in Section 2.
@@
$\gamma^{i}_{t}(h)$ $\approx\gamma^{i}_{t}(0)+h\dot{\gamma}^{i}_{t}(0)\Rightarrow\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta g^{ij}\partial_{j}\mathcal{L}.$ (19)
where the learning rate $\eta=h\lambda$. This approximation corresponds to the naive Quantum Natural Gradient update rule, utilizing only the first-order information. For a more precise approximation, we can incorporate second-order information from the geodesic equation:
$\gamma^{i}_{t}(h)$ $\approx\gamma^{i}_{t}(0)+h\dot{\gamma}^{i}_{t}(0)+\frac{1}{2}h^{2}\ddot{\gamma}^{i}_{t}(0),$ (20)
where, according to the geodesic equation (3):
$\ddot{\gamma}^{i}_{t}(0)$ $=-\frac{1}{2}\Gamma^{i}_{lm}\dot{\gamma}^{l}_{t}(0)\dot{\gamma}^{m}_{t}(0).$ (21)
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}+h\dot{\gamma}^{i}_{t}(0)-\frac{1}{2}h^{2}\Gamma^{i}_{lm}\dot{\gamma}^{l}_{t}(0)\dot{\gamma}^{m}_{t}(0),$ (22)
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta g^{ij}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})-\frac{b}{2}\Gamma^{i}_{lm}(\bm{\theta}_{t})\left(g^{lj}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})\right)\left(g^{mj}(\bm{\theta}_{t})\partial_{j}\mathcal{L}(\bm{\theta}_{t})\right).$ (23)
@@
5 Computing the Christoffel Symbols with the Parameter-Shift Rule
The Christoffel symbols, $\Gamma^{i}_{jk}$, are essential components in the update rule (23), as they capture the curvature of the parameter space defined by the Fubini-Study metric. Traditionally, these symbols are computed by differentiating the Fubini-Study metric (15), which involves first calculating the metric tensor and then deriving the Christoffel symbols through classical differentiation. This approach relies heavily on classical post-processing and does not leverage the quantum device directly.
To enable direct estimation of the Christoffel symbols on a quantum device, we reformulate their computation using the parameter-shift rule specifically adapted to the quantum circuit of interest. This method allows the Christoffel symbols to be directly accessed through quantum measurements, bypassing the need for classical differentiation and making the process more efficient and directly linked to the quantum state preparation. For a detailed overview of the parameter-shift rule and its application in deriving the Fubini-Study metric, please refer to the appendix and the related work in [26].

### v3

@@ I Introduction
The Quantum Natural Gradient (QNG) algorithm [6], a generalization of the natural gradient [7], is an advanced optimization technique that enhances the optimization process by incorporating the geometry of the parameter space into the update rules. Unlike standard gradient descent, which assumes a flat parameter space, QNG leverages a Riemannian metric defined by the Fubini–Study metric or, more generally, by the Quantum Fisher Information Matrix (QFIM) [8], capturing the infinitesimal changes in distances between quantum states and aligning parameter updates with the natural curvature of the quantum state manifold. A key factor in QNG’s superior performance, as shown in [9], is its ability to identify regions of high negative curvature early in the optimization process, significantly accelerating convergence. These regions play a crucial role in guiding the optimization along paths that lead to faster descent.
@@
This manuscript presents a comprehensive approach to incorporating geodesic corrections into the QNG for applications in variational quantum algorithms. In Section II, we provide an overview of differential geometry and geodesic equations, establishing the foundational mathematical context. Section III delves into optimizing idealized variational quantum circuits using Quantum Natural Gradient Descent (QNG), emphasizing its effectiveness in quantum optimization. Section IV introduces higher-order integrators and derives the update rule for the Quantum Natural Gradient with Geodesic Correction (QNGGC). As the update rule relies on the Christoffel symbols of the second kind, Section V is dedicated to efficiently computing these symbols using the parameter-shift rule, enabling direct measurements from quantum circuits. In Section VI, we apply the derived update rule to various examples: Examples 1 and 2 involve analytical calculations for numerical simulations, while Example 3 utilizes quantum software, specifically Qiskit [30], for practical simulations. Finally, in Section VII, we provide an outlook on potential extensions and future research directions, highlighting the integration of geodesic corrections into quantum optimization frameworks.
II Differential Geometry and Geodesic Equations
@@
A manifold is a topological space that locally resembles Euclidean space and supports a consistent coordinate system. More formally, an $n$-dimensional manifold $\mathcal{M}$ is a set equipped with a collection of coordinate charts $\{(U_{i},\varphi_{i})\}_{i\in I}$, where $I$ is an index set and each $U_{i}\subset\mathcal{M}$ is an open subset , and $\varphi_{i}:U_{i}\rightarrow\mathbb{R}^{n}$ is a homeomorphism, meaning that $U_{i}$ is locally similar to $\mathbb{R}^{n}$. For $\mathcal{M}$ to be a differentiable manifold, the transition maps between overlapping charts, $\varphi_{j}\circ\varphi_{i}^{-1}:\varphi_{i}(U_{i}\cap U_{j})\rightarrow\varphi_{j}(U_{i}\cap U_{j})$, must be smooth (infinitely differentiable). This smooth structure allows us to perform calculus on the manifold.
@@
where $g_{ij}=g_{ji}$, and the indices $i,j$ run from $1$ to $n$, with $n$ being the dimensionality of the manifold $\mathcal{M}$. The metric allows for measuring lengths and angles on the manifold. The length of a smooth curve $\gamma:[a,b]\rightarrow\mathcal{M}$ is given by:
@@
The Christoffel symbols define the covariant derivative, which maps tensor fields to other tensor fields, adapting to the curvature of the manifold. Specifically, the covariant derivative of a vector field $V^{j}$ is defined as:
$\nabla_{i}V^{j}=\partial_{i}V^{j}+\Gamma^{j}_{ik}V^{k},$ (5)
This expression generalizes the notion of directional derivatives to curved manifolds and ensures invariance under changes of coordinates.
The exponential map is a tool in Riemannian geometry that relates the tangent space at a point to the manifold. Given a tangent vector $v\in T_{p}\mathcal{M}$, the exponential map $\text{Exp}_{p}$ maps $v$ to a point on the manifold reached by traveling along the geodesic starting at $p$ with initial velocity $v$. Formally,
@@
where $\gamma(\tau)$ is the geodesic satisfying the initial conditions $\gamma(0)=p$ and $\dot{\gamma}(0)=\left.\frac{d\gamma(\tau)}{d\tau}\right|_{\tau=0}=v$. By rescaling the parameter $v$ by a factor of $\epsilon$, the following relation holds:
$\text{Exp}_{p}(\epsilon v)=\gamma(\epsilon).$ (7)
To approximate $\text{Exp}_{p}(\epsilon v)$ locally, we expand $\gamma(\epsilon)$ for small $\epsilon$ using a Taylor series:
$\text{Exp}_{p}(\epsilon v)\approx\gamma(0)+\epsilon\dot{\gamma}(0)+\frac{\epsilon^{2}}{2}\ddot{\gamma}(0)+\mathcal{O}(\epsilon^{3}).$ (8)
This expansion provides insight into the behavior of the geodesic near the starting point $p$. The first-order term, $\gamma(0)+\epsilon\dot{\gamma}(0)$, corresponds to the initial position and velocity, providing a linear approximation of the geodesic. The second-order term, $\frac{\epsilon^{2}}{2}\ddot{\gamma}(0)$, accounts for curvature effects, refining the approximation by incorporating the geodesic’s acceleration. Since $\ddot{\gamma}(0)$ is given by evaluating the geodesic equation (3) at $\tau=0$, it encodes the local curvature effects at the starting point.
@@
III Optimizing Idealized Variational Quantum Circuits with Quantum Natural Gradient
@@
where $V_{j}$ are fixed unitary operators, and $U_{j}(\theta_{j})$ are parameterized gates. Here, $\bm{\theta}=(\theta_{1},\dots,\theta_{l})$ is the vector of parameters, and the index $l$ denotes the total number of parameterized gates in the circuit. Each $U_{j}(\theta_{j})$ is defined as:
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta\,\bm{\nabla}\mathcal{L}(\bm{\theta}_{t}),$ (12)
where $\bm{\nabla}:=(\partial_{1},\dots,\partial_{l})=\left(\frac{\partial}{\partial\theta^{1}},\dots,\frac{\partial}{\partial\theta^{l}}\right)$ denotes the gradient operator with respect to the parameter vector $\bm{\theta}=(\theta_{1},\dots,\theta_{l})$.
Equation (12) can be interpreted as an approximation of the solution to an ordinary differential equation (ODE) using the Euler method:
$\dot{\bm{\theta}}$ $=-\lambda\,\bm{\nabla}\mathcal{L}(\bm{\theta}),$ (13)
where $\eta=\epsilon\lambda$ is the learning rate, with $\lambda$ being a time scale constant that affects the speed but not the trajectory of the system, and $\epsilon$ is the step size.
However, this ODE is not invariant under reparameterizations of the parameters $\bm{\theta}$. For instance, if we rescale the parameters $\bm{\theta}\rightarrow 2\bm{\theta}$, the gradient $\bm{\nabla}\mathcal{L}$ would scale as $\frac{1}{2}\bm{\nabla}\mathcal{L}$, leading to inconsistencies in the optimization process.
The core of this issue lies in the differential geometric nature of the gradient. The parameter update $\dot{\bm{\theta}}$ transforms as a vector in the tangent space $T_{\bm{\theta}}\mathcal{M}$, while the gradient $\bm{\nabla}\mathcal{L}$ is a covector (or 1-form) in the cotangent space $T^{*}_{\bm{\theta}}\mathcal{M}$. Since the ODE in Eq. (12) attempts to relate objects in different spaces with distinct transformation rules, it is not an invariant relation.
QNG alleviates this issue by approximately solving an invariant ODE. The key idea is to apply the inverse metric $\mathbf{g}^{-1}$ to the gradient vector, yielding the natural gradient $\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta})$. The resulting ODE becomes:
$\dot{\bm{\theta}}=-\lambda\,\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta}),$ (14)
@@
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta}_{t}),$ (15)
remains consistent across different parameter spaces. The metric tensor $g_{ij}$ is recognized as the Fubini-Study metric, derived from the real part of the Quantum Geometric Tensor [32]:
@@
IV Higher-order Integrators and Geodesic Correction
@@
$\bm{\gamma}_{t}(\epsilon)$ $=\bm{\theta}_{t+1}=\text{Exp}_{\bm{\theta}_{t}}\left(-\epsilon\lambda\,\mathbf{g}^{-1}(\bm{\theta}_{t})\,\bm{\nabla}\mathcal{L}(\bm{\theta}_{t})\right),$ (17)
@@
$\bm{\gamma}_{t}(0)$ $=\bm{\theta}_{t},$ (18)
$\dot{\bm{\gamma}}_{t}(0)$ $=-\lambda\,\mathbf{g}^{-1}(\bm{\theta}_{t})\,\bm{\nabla}\mathcal{L}(\bm{\theta}_{t}).$ (19)
Here, the exponential map Exp translates the current parameter $\bm{\theta}_{t}$ along the geodesic defined by the natural gradient. This approach is effective because it preserves invariance properties under reparameterization, owing to the characteristics of the exponential map. However, directly computing the exponential map is challenging, as it requires solving the geodesic equation exactly. To approximate this computation, we explore using the first and second-order derivatives as detailed in Section II.
@@
$\bm{\gamma}_{t}(\epsilon)\approx\bm{\gamma}_{t}(0)+\epsilon\dot{\bm{\gamma}}_{t}(0)\Rightarrow\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\,\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta}_{t}),$ (20)
where the learning rate $\eta=\epsilon\lambda$. This approximation corresponds to the naive Quantum Natural Gradient update rule, utilizing only the first-order information. For a more precise approximation, we can incorporate second-order information from the geodesic equation, we have:
$\bm{\gamma}_{t}(\epsilon)\approx\bm{\gamma}_{t}(0)+\epsilon\dot{\bm{\gamma}}_{t}(0)+\frac{1}{2}\epsilon^{2}\ddot{\bm{\gamma}}_{t}(0).$ (21)
Evaluating the geodesic equation (3) at $\tau=0$ gives:
$\ddot{\bm{\gamma}}_{t}(0)=-\,\bm{\Gamma}(\bm{\theta}_{t})\left[\dot{\bm{\gamma}}_{t}(0),\dot{\bm{\gamma}}_{t}(0)\right].$ (22)
Here, $\bm{\Gamma}[\cdot,\cdot]$ denotes the action of the connection (Christoffel symbols) as a bilinear map on tangent vectors; in components, it corresponds to $\Gamma^{i}_{lm}\dot{\gamma}^{l}\dot{\gamma}^{m}$.
@@
$\bm{\theta}_{t+1}=\bm{\theta}_{t}+\epsilon\dot{\bm{\gamma}}_{t}(0)-\frac{1}{2}\epsilon^{2}\,\bm{\Gamma}(\bm{\theta}_{t})\left[\dot{\bm{\gamma}}_{t}(0),\dot{\bm{\gamma}}_{t}(0)\right].$ (23)
@@
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\,\mathbf{g}^{-1}(\bm{\theta}_{t})\bm{\nabla}\mathcal{L}(\bm{\theta}_{t})-\frac{b}{2}\,\bm{\Gamma}(\bm{\theta}_{t})\left[\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta}_{t}),\,\mathbf{g}^{-1}\bm{\nabla}\mathcal{L}(\bm{\theta}_{t})\right].$ (24)
This is equivalent, in components, to the expression:
$\theta^{\mu}_{t+1}=\theta^{\mu}_{t}-\eta\,g^{ij}(\theta^{\mu}_{t})\,\partial_{j}\mathcal{L}(\theta^{\mu}_{t})-\frac{b}{2}\,\Gamma^{i}_{lm}(\theta^{\mu}_{t})\left(g^{lj}(\theta^{\mu}_{t})\,\partial_{j}\mathcal{L}(\theta^{\mu}_{t})\right)\left(g^{mj}(\theta^{\mu}_{t})\,\partial_{j}\mathcal{L}(\theta^{\mu}_{t})\right).$ (25)
where $\theta^{\mu}$ denotes the components of the parameter vector $\bm{\theta}$ in a local coordinate, and $\partial_{j}:=\frac{\partial}{\partial\theta^{j}}$.
@@
V Computing the Christoffel Symbols with the Parameter-Shift Rule
The Christoffel symbols are essential components in the update rule (24), as they capture the curvature of the parameter space defined by the Fubini-Study metric. Traditionally, the Christoffel symbols are computed by differentiating the Fubini-Study metric (16), which involves first calculating the metric tensor and then deriving the Christoffel symbols through classical differentiation. This approach relies heavily on classical post-processing and does not fully utilize information available from parameterized quantum states.
To enable direct estimation of the Christoffel symbols, we reformulate their computation using the parameter-shift rule specifically adapted to parameterized quantum circuits. This method allows the Christoffel symbols to be directly estimated from measurement outcomes, bypassing the need for classical differentiation and making the process more efficient and directly linked to quantum state preparation. For a detailed overview of the parameter-shift rule and its application in deriving the Fubini-Study metric, please refer to the appendix and the related work in [26].